Tennis court length is 3ft less than three times its width, area is 2106 ft squared
What is length
What is width
I started with quadratic equation
ax (squared) + bx + c = 0
A= lxw
L= 3w -3
L-3 x 3w= 2106ft squared
But got lost with the area squared
Thanks in advance for your help
Dianne

Answer by stanbon(75887)

Tennis court length is 3ft less than three times its width,
area is 2106 ft squared
-----
Let width be "w":
Then length is "3w-3"
-----
Area = length * width
2106 = (3w-3)w
3w^2 -3w - 2106 = 0
w^2 -w - 702 = 0
----
w = 27 ft (width)
length = 3x-3 = 78 ft (length)
==================
Cheers,
Stan H.
==================